OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

i think u have to substitute sin x with y 7y^2 - 14y + 2 = -5 add 5 to both sides 7y^2 - 14y + 7 = 0 divide everything by 7 y^2 - 2 y + 1 = 0 (y - 1)(y - 1) = 0 so the solution is y =1 y = sin (x) = 1 when is sin(x) = 1?

OpenStudy (anonymous):

Surdawl did exactly what I was about to type!

OpenStudy (anonymous):

so there is only one solution?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

seems right to me.

OpenStudy (anonymous):

thanks so much!